# Upper Approximation Bounds for Neural Oscillators

### v1

@@ B.3 Proof of Theorem 2
First, the first part on the right side of Eq. (98) is considered. According to the assumptions in Theorem 2, since $g(\cdot)$ is Lipschitz continuous over $\mathbb{R}^{2r+p}$, the solutions ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$ to Eq. (29) driven by all $\mathbf{u}(t)\in K$, with the initial conditions ${\hat{\mathbf{x}}}(0)={\hat{\mathbf{x}}}^{\prime}(0)=\mathbf{0}$, are unique and they can be bounded by two constants $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$, respectively, that is $\left|{\hat{x}_{j}}(t)\right|\leq B_{\hat{\mathbf{x}}}$ and $\left|{\hat{x}_{j}^{\prime}}(t)\right|\leq B_{\hat{\mathbf{x}}^{\prime}}$ for all $\mathbf{u}(t)\in K$, all $t\in[0,T]$, and j=1,2,,rj=1,2,...,r. Since the dynamical system in Eq. (29) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}$, where $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, and $B_{\beta{g}}$ is the stability bound of the dynamical system in Eq. (29), which is calculated by Eq. (25), if the difference between the MLP ${\it{\Gamma}}(\cdot)$ in Eq. (1) and $g(\cdot)$ in Eq. (29) over $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ is bounded by
@@
$\begin{array}[]{l}\left|{{\bf{x}}(t)-{\hat{\bf x}}(t)}\right|+\left|{{\bf{x^{\prime}}}(t)-{\hat{\bf x}^{\prime}}(t)}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x^{\prime}}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x^{\prime}}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{{\beta_{g}}}}\leq\alpha{B_{{\beta_{g}}}},\end{array}$ (100)
@@
Since each $g_{j}(\cdot)$ of $g(\cdot)$ has arbitrary finite-order derivatives over $\mathbb{R}^{2r+p}$, there exists a corresponding ${\tilde{g}}_{j}(\cdot)$ equaling $g_{j}(\cdot)$ over $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ and the first moment of the Fourier transform $\mathcal{H}_{{\tilde{g}}_{j}}\left(\boldsymbol{\upomega}\right)$ of ${\tilde{g}}_{j}(\cdot)$ is finite (Barron, 1993), that is
$\int_{{\mathbb{R}^{2r+p}}}{\left|\boldsymbol{\upomega}\right|_{2}}\left|{{\mathcal{H}_{{{\tilde{g}}_{j}}}}(\boldsymbol{\upomega})}\right|{\rm{d}}\boldsymbol{\upomega}<+\infty,\;\;j=1,2,\ldots,r,$ (101)
where $\boldsymbol{\upomega}=\left[\omega_{1},\omega_{2},...,\omega_{2r+p}\right]^{\top}$. Based on Theorem 2.2 in Yukich et al. (1995) and the fact that four ReLU activation functions can form a total variation activation function that has a bounded support, e.g., $\left[\sigma_{\mathrm{ReLU}}(x)-\sigma_{\mathrm{ReLU}}(x-1)\right]-\left[\sigma_{\mathrm{ReLU}}(x-2)-\sigma_{\mathrm{ReLU}}(x-3)\right]$, for an arbitrary $\varepsilon_{1}$ satisfying $0<\varepsilon_{1}<\alpha/r$, there exist $r$ one-hidden-layer MLPs ${\it{\Gamma}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, j=1,2,,rj=1,2,...,r, the widths of whose input and output layers are respectively $2r+p$ and 1, and the width of the hidden layer for each ${\it{\Gamma}}_{j}(\cdot)$ is at most
@@
Then, from Eqs. (99), (100), and (102), for an arbitrary $\varepsilon_{1}$ satisfying $0<\varepsilon_{1}<\alpha/r$, there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ consisting of above ${\it{\Gamma}}_{1}(\cdot),{\it{\Gamma}}_{2}(\cdot),...,{\it{\Gamma}}_{r}(\cdot)$, the widths of the input and output layers of ${\it{\Gamma}}(\cdot)$ are respectively $2r+p$ and $r$, and the width of its hidden layer is at most
@@
$\begin{array}[]{l}\left|{{\bf{x}}(t)-{\hat{\bf x}}(t)}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{{\beta_{g}}}}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq{B_{{\beta_{g}}}}\sqrt{\sum\limits_{j=1}^{r}{\mathop{\max}\limits_{\tau\in[0,T]}\left\{{{{\it{\Gamma}}_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-{g_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right\}^{2}}}\leq{B_{{\beta_{g}}}}\sqrt{r}{\varepsilon_{1}}\end{array}$ (105)
@@
$\begin{array}[]{l}\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}\\ \;\;\;\;\leq\sqrt{\sum\limits_{j=1}^{r}{\mathop{\max}\limits_{\tau\in[0,T]}\left\{{{{\it{\Gamma}}_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-{g_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right\}^{2}}}\leq\sqrt{r}{\varepsilon_{1}}\leq\alpha{r^{-0.5}}\leq\alpha.\end{array}$ (106)
The first term on the right side of Eq. (98) can be bounded by
@@
for all $t\in[0,T]$, where $L_{h}$ is the Lipschitz constant of the function $h(\cdot)$.
Subsequently, the second term on the right side of Eq. (98) is taken into account. Using the same method based on the Barron class (Barron, 1993), since $h(\cdot)$ has arbitrary finite-order derivatives over $\mathbb{R}^{r}$, for each $j^{\text{th}}$ element $h_{j}(\cdot)$ of $h(\cdot)$, j=1,2,,qj=1,2,...,q, there exists a corresponding ${\tilde{h}}_{j}(\cdot)$ equaling $h_{j}(\cdot)$ over $\Omega\left(B_{\mathbf{x}}^{r}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}$, the first moment of the Fourier transform $\mathcal{H}_{{\tilde{h}}_{j}}\left(\boldsymbol{\upomega}\right)$ of ${\tilde{h}}_{j}(\cdot)$ is finite (Barron, 1993). Based on Theorem 2.2 in Yukich et al. (1995) and the fact that four ReLU activation functions can form a total variation activation function that has a bounded support, for an arbitrary positive $\varepsilon_{2}$, there exists q one-hidden-layer MLPs ${\it{\Pi}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the parameters in each ${\it{\Pi}}_{j}(\cdot)$ with respect to $\mathbf{u}(0)$ and $t$ are zero, the widths of the input and output layers of each ${\it{\Pi}}_{j}(\cdot)$ are respectively $r+p+1$ and 1, the width of the hidden layer for each ${\it{\Pi}}_{j}(\cdot)$ is at most
@@ Appendix C Explanation of the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$
Assume that $\mathbf{u}(0)=\mathbf{0}$ holds for all $\mathbf{u}(t)\in K$, then, the continuous mapping $\Psi\left({\bf{x}},{\bf{0}},t\right)$ in Lemma 3 is now defined as $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)=\Psi\left[{\mathbf{x}}(t),{\bf{0}},t\right]=\Phi[{\hat{\mathbf{u}}}_{t}(\tau)](t)$, where the approximate ${\hat{\mathbf{u}}}_{t}(\tau)$ of $\mathbf{u}(t)$ is calculated as $\hat{\mathbf{u}}_{t}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}(t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}$ and $\mathbf{x}_{n}(t)=\mathcal{L}_{t}\mathbf{u}(\omega_{n})$ in Eq. (55). For $\forall\mathbf{u}_{1}(t)\in K$ with a positive $\Delta t$, its shifted version $\mathbf{u}_{2}(t)$ is
@@
For $\forall t_{1}\geq 0$ and $t_{2}=t_{1}+\Delta t\leq T$, it is satisfied $\mathbf{x}_{1,n}(t_{1})=\mathcal{L}_{t_{1}}\mathbf{u}_{1}(\omega_{n})=\mathcal{L}_{t_{2}}\mathbf{u}_{2}(\omega_{n})=\mathbf{x}_{2,n}(t_{2})$, which is denoted as $\mathbf{x}_{1,n}(t_{1})=\mathbf{x}_{2,n}(t_{2})=\mathbf{x}_{n}$. Then, the approximates ${\hat{\mathbf{u}}_{1,t_{1}}}(\tau)$ and ${\hat{\mathbf{u}}_{2,t_{2}}}(\tau)$ of $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ are respectively ${\hat{\mathbf{u}}_{1,t_{1}}}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]}$ and ${\hat{\mathbf{u}}_{2,t_{2}}}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]}$. The related $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{1}\right)$ is $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{1}\right)$ $\Phi\left[\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]}\right](t_{1})$ and $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)$ is $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)=\Phi\left[\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]}\right](t_{2})$. Although $\Phi$ is assumed to be time-autonomous, thereby $\Phi\left[\mathbf{u}_{1}(\tau)\right](t_{1})=\Phi\left[\mathbf{u}_{2}(\tau)\right](t_{2})$, $\Psi({\mathbf{x}}_{n},{\mathbf{0}},$ $t_{1})$ does not equal $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)$. This is because the valid range of $\tau$ in $\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]$ is $0,t_{1}]$ but that of $\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]$ is $0,t_{2}]$. Thus, the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$ cannot be ignored even if $\Phi$ is time-autonomous.
An alternative approximate ${\hat{\mathbf{u}}_{2,t_{2},\Delta t}}(\tau)$ of $\mathbf{u}_{2}(t)$ is
@@
It is satisfied $\Phi\left[\hat{\mathbf{u}}_{1,t_{1}}(\tau)\right](t_{1})=\Phi\left[\hat{\mathbf{u}}_{2,t_{2},\Delta t}(\tau)\right](t_{2})$. However, the approximate in Eq. (113) cannot apply to the input functions with non-zero values over $0,\Delta t)$, and thus it is not applicable to all functions in K.

### v2

@@ B.3 Proof of Theorem 2
First, the first part on the right side of Eq. (103) is considered. According to the assumptions in Theorem 2, for every compact domain $\Omega\subset{\mathbb{R}}^{2r+p}$, each $g_{j}(\cdot)$ of $g(\cdot)$ has a Barron function extension $\tilde{g}_{j,\Omega}(\cdot)=g_{j}(\cdot)$ on $\Omega$, and the spectral Barron norms $\|{g}_{j}\|_{\mathcal{B},\Omega}=\int_{{\mathbb{R}^{2r+p}}}{\left(1+\left|\boldsymbol{\upomega}\right|\right)}\left|{{\mathcal{H}_{{{\tilde{g}}_{j,\Omega}}}}(\boldsymbol{\upomega})}\right|{\rm{d}}\boldsymbol{\upomega}$ (Siegel and Xu, 2023) of all extensions are uniformly bounded, where $\mathcal{H}_{{\tilde{g}}_{j,\Omega}}\left(\boldsymbol{\upomega}\right)$ is the Fourier transform of ${\tilde{g}}_{j,\Omega}(\cdot)$ and $\boldsymbol{\upomega}=[\omega_{1},\omega_{2},\dots,\omega_{2r+p}]^{\top}$ is a radian frequency vector. Thus, $g(\cdot)$ is bounded and continuously differentiable with bounded derivatives over $\mathbb{R}^{2r+p}$ (Barron, 1993). The solutions ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$ to Eq. (29) driven by all $\mathbf{u}(t)\in K$, with the initial conditions ${\hat{\mathbf{x}}}(0)={\hat{\mathbf{x}}}^{\prime}(0)=\mathbf{0}$, are unique and they can be bounded by two constants $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$, respectively, that is $\left|{\hat{x}_{j}}(t)\right|\leq B_{\hat{\mathbf{x}}}$ and $\left|{\hat{x}_{j}^{\prime}}(t)\right|\leq B_{\hat{\mathbf{x}}^{\prime}}$ for all $\mathbf{u}(t)\in K$, all $t\in[0,T]$, and j=1,2,,rj=1,2,...,r. Since the dynamical system in Eq. (29) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}$, where $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, and $B_{\beta{g}}$ is the stability bound of the dynamical system in Eq. (29), which is calculated by Eq. (25), if the difference between the MLP ${\it{\Gamma}}(\cdot)$ in Eq. (1) and $g(\cdot)$ in Eq. (29) over $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ is bounded by
@@
$\left|{{\bf{x}}(t)-{\hat{\bf x}}(t)}\right|+\left|{{\bf{x^{\prime}}}(t)-{\hat{\bf x}^{\prime}}(t)}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x^{\prime}}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x^{\prime}}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{{\beta_{g}}}}\leq\alpha{B_{{\beta_{g}}}},$ (105)
@@
Since each $g_{j}(\cdot)$ of $g(\cdot)$ has a Barron function extension $\tilde{g}_{j,\Omega}(\cdot)=g_{j}(\cdot)$ on $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ with a finite spectral Barron norm, based on Theorem 2.2 in Yukich et al. (1995) and the fact that four ReLU activation functions can form a total variation activation function that has a bounded support, e.g., $\left[\sigma_{\mathrm{ReLU}}(x)-\sigma_{\mathrm{ReLU}}(x-1)\right]-\left[\sigma_{\mathrm{ReLU}}(x-2)-\sigma_{\mathrm{ReLU}}(x-3)\right]$, for an arbitrary $\varepsilon_{1}$ satisfying $0<\varepsilon_{1}<\alpha/r$, there exist $r$ one-hidden-layer MLPs ${\it{\Gamma}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, j=1,2,,rj=1,2,...,r, the widths of whose input and output layers are respectively $2r+p$ and 1, and the width of the hidden layer for each ${\it{\Gamma}}_{j}(\cdot)$ is at most
@@
Then, from Eqs. (104)-(106), for an arbitrary $\varepsilon_{1}$ satisfying $0<\varepsilon_{1}<\alpha/r$, there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ consisting of above ${\it{\Gamma}}_{1}(\cdot),{\it{\Gamma}}_{2}(\cdot),...,{\it{\Gamma}}_{r}(\cdot)$, the widths of the input and output layers of ${\it{\Gamma}}(\cdot)$ are respectively $2r+p$ and $r$, and the width of its hidden layer is at most
@@
$\left|{{\bf{x}}(t)-{\hat{\bf x}}(t)}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{{\beta_{g}}}}$ (109)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq{B_{{\beta_{g}}}}\sqrt{\sum\limits_{j=1}^{r}{\mathop{\max}\limits_{\tau\in[0,T]}\left\{{{{\it{\Gamma}}_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-{g_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right\}^{2}}}\leq{B_{{\beta_{g}}}}\sqrt{r}{\varepsilon_{1}}$
@@
$\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}$ (110)
$\;\;\;\;\leq\sqrt{\sum\limits_{j=1}^{r}{\mathop{\max}\limits_{\tau\in[0,T]}\left\{{{{\it{\Gamma}}_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-{g_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right\}^{2}}}\leq\sqrt{r}{\varepsilon_{1}}\leq\alpha{r^{-0.5}}\leq\alpha.$
Since each $h_{j}(\cdot)$ of $h(\cdot)$, j=1,2,,qj=1,2,...,q, has a Barron function extension ${\tilde{h}}_{j,\Omega}(\cdot)=h_{j}(\cdot)$ on $\Omega\left(B_{\mathbf{x}}^{r}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}$, and the spectral Barron norm $\|{h}_{j}\|_{\mathcal{B},\Omega}=\int_{{\mathbb{R}^{r}}}{\left(1+\left|\boldsymbol{\upomega}\right|\right)}\left|{{\mathcal{H}_{{{\tilde{h}}_{j,\Omega}}}}(\boldsymbol{\upomega})}\right|{\rm{d}}\boldsymbol{\upomega}$ of each ${\tilde{h}}_{j,\Omega}(\cdot)$ is finite, $h(\cdot)$ is Lipschitz continuous with a Lipschitz constant $L_{h}$. The first term on the right side of Eq. (103) can be bounded by
@@
for all $t\in[0,T]$.
Subsequently, the second term on the right side of Eq. (103) is taken into account. Using the same method based on the Barron class (Barron, 1993), Theorem 2.2 in Yukich et al. (1995), and the fact that four ReLU activation functions can form a total variation activation function that has a bounded support, for an arbitrary positive $\varepsilon_{2}$, there exists q one-hidden-layer MLPs ${\it{\Pi}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the parameters in each ${\it{\Pi}}_{j}(\cdot)$ with respect to $\mathbf{u}(0)$ and $t$ are zero, the widths of the input and output layers of each ${\it{\Pi}}_{j}(\cdot)$ are respectively $r+p+1$ and 1, the width of the hidden layer for each ${\it{\Pi}}_{j}(\cdot)$ is at most
@@ Appendix C Explanation of the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$
Assume that $\mathbf{u}(0)=\mathbf{0}$ holds for all $\mathbf{u}(t)\in K$, then, the continuous mapping $\Psi\left({\bf{x}},{\bf{0}},t\right)$ in Lemma 3 is now defined as $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)=\Psi\left[{\mathbf{x}}(t),{\bf{0}},t\right]=\Phi[{\hat{\mathbf{u}}}_{t}(\tau)](t)$, where the approximation ${\hat{\mathbf{u}}}_{t}(\tau)$ of $\mathbf{u}(t)$ is calculated as $\hat{\mathbf{u}}_{t}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}(t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}$ and $\mathbf{x}_{n}(t)=\mathcal{L}_{t}\mathbf{u}(\omega_{n})$ in Eq. (60). For $\forall\mathbf{u}_{1}(t)\in K$ with a positive $\Delta t$, its shifted version $\mathbf{u}_{2}(t)$ is
@@
For $\forall t_{1}\geq 0$ and $t_{2}=t_{1}+\Delta t\leq T$, it is satisfied $\mathbf{x}_{1,n}(t_{1})=\mathcal{L}_{t_{1}}\mathbf{u}_{1}(\omega_{n})=\mathcal{L}_{t_{2}}\mathbf{u}_{2}(\omega_{n})=\mathbf{x}_{2,n}(t_{2})$, which is denoted as $\mathbf{x}_{1,n}(t_{1})=\mathbf{x}_{2,n}(t_{2})=\mathbf{x}_{n}$. Then, the approximations ${\hat{\mathbf{u}}_{1,t_{1}}}(\tau)$ and ${\hat{\mathbf{u}}_{2,t_{2}}}(\tau)$ of $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ are respectively ${\hat{\mathbf{u}}_{1,t_{1}}}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]}$ and ${\hat{\mathbf{u}}_{2,t_{2}}}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]}$. The related $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{1}\right)$ is $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{1}\right)$ $\Phi\left[\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]}\right](t_{1})$ and $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)$ is $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)=\Phi\left[\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]}\right](t_{2})$. Although $\Phi$ is assumed to be time-autonomous, thereby $\Phi\left[\mathbf{u}_{1}(\tau)\right](t_{1})=\Phi\left[\mathbf{u}_{2}(\tau)\right](t_{2})$, $\Psi({\mathbf{x}}_{n},{\mathbf{0}},$ $t_{1})$ does not equal $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)$. This is because the valid range of $\tau$ in $\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]$ is $0,t_{1}]$ but that of $\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]$ is $0,t_{2}]$. Thus, the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$ cannot be ignored even if $\Phi$ is time-autonomous.
An alternative approximation ${\hat{\mathbf{u}}_{2,t_{2},\Delta t}}(\tau)$ of $\mathbf{u}_{2}(t)$ is
@@
It is satisfied $\Phi\left[\hat{\mathbf{u}}_{1,t_{1}}(\tau)\right](t_{1})=\Phi\left[\hat{\mathbf{u}}_{2,t_{2},\Delta t}(\tau)\right](t_{2})$. However, the approximation in Eq. (117) cannot apply to the input functions with non-zero values over $0,\Delta t)$, and thus it is not applicable to all functions in K.